OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of –3, 5 of multiplicity 2. Explain how you would verify the zeros of f(x).

OpenStudy (anonymous):

if it has zeros at those points, then it must have factors (x+3)(x-5) as if you set this equal to = x +3 = 0 and x - 5 = 0 you will see x = -3, 5 (which was the original zeros) and multiplicity of 2 just means it is scaled by 2 thus 2(x+3)(x-5) = f(x)

OpenStudy (anonymous):

thank you!!!!!